OpenStudy (anonymous):

Are 20, 25 and 30 an example of a Pythagorean triplet? Yes or No

jimthompson5910 (jim_thompson5910):

If you can show that 20^2+25^2 = 30^2, then you will have shown that 20,25,30 is a pythagorean triple.

Directrix (directrix):

No. 20 = 5 *4 25 = 5*5 30 = 5*6 (4, 5, 6) is not a Pythagorean Triple -------------

jimthompson5910 (jim_thompson5910):

Right off the bat however, the units digit of 20^2 will be 0 and the units digit of 25^2 will be 5. These units digits add to 0+5 = 5, which is NOT the units digit of 30^2 (which is 0)

OpenStudy (anonymous):

No If they are a Pythagorean triple, then Pythagoras's Theorem applies: 30² = 20² + 25² But 30² = 900 And 20² + 25² = 400 + 625 = 1025. Not 900. So not a Pythagorean triple.

OpenStudy (anonymous):

thanks for the list

OpenStudy (anonymous):

npz

OpenStudy (anonymous):

so bst answer go's toooooooooooooooooooooooo.......................